It has been estimated that a man’s heart is about 33 cm below his brain. Considering the density 

of human blood to be 1.1 x 103 kg/m3

, what pressure must the heart produce to supply blood to 

the brain?


1
Expert's answer
2021-05-13T11:02:16-0400

"\text{initial data:}"

"\rho = 1.1*10^3\ \frac{kg}{m^3}"

"h = 33cm = 0.33m"

"g = 9.8 \frac{m}{s^2}"

"\text{Let us find the pressure inside a column of blood with a height of h:}"

"p =\rho gh=0.33*1.1*10^3*9.8=3557.4 \ Pa"

"\text{the pressure p found must be produced by the heart }"

"\text{o raise the blood to the height h of the brain}"

"\text{Answer: }3557.4\ Pa"